# Identifying and Analyzing Performance-Critical Tokens in Large Language Models

### v3

@@ Appendix I Analysis of the Next-line Tokens
The full results on all the six datasets are shown in Table 29 and Table 30. Most of the results align with our descriptions in Section C, where the performance-critical tokens should be utilized together to provide the best performance and that removing some of them would cause performance degeneration, demonstrated by the performance decrease from Table 29, or instability issues, shown by Table 30.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||
with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | ||
OpenLlama 3B | temp with | 56.5 | 47.4 | 86.7 | 83.7 | 27.1 | 26.5 | 62.2 | 59.8 | 56.4 | 56.0 | 52.3 | 56.1 |
50.3 | 47.1 | 85.7 | 84.4 | 28.9 | 24.4 | 57.7 | 57.7 | 56.5 | 56.1 | 53.2 | 55.2 | ||
34.6 | 32.7 | 86.9 | 82.3 | 28.2 | 31.2 | 55.5 | 54.1 | 58.3 | 59.2 | 55.4 | 58.3 | ||
Llama 7B | temp with | 70.8 | 57.3 | 90.2 | 87.1 | 58.4 | 46.7 | 66.2 | 63.8 | 66.3 | 59.5 | 73.5 | 69.6 |
62.7 | 55.1 | 91.6 | 87.1 | 52.8 | 43.3 | 61.6 | 61.8 | 58.9 | 56.5 | 59.2 | 55.7 | ||
50.8 | 48.6 | 84.9 | 82.8 | 46.0 | 50.2 | 57.9 | 55.2 | 56.7 | 56.7 | 66.2 | 64.5 | ||
Llama 13B | temp with | 80.0 | 76.2 | 92.3 | 89.1 | 58.6 | 54.0 | 76.9 | 71.4 | 68.5 | 59.8 | 47.7 | 35.0 |
79.9 | 76.3 | 91.5 | 88.9 | 55.1 | 47.8 | 75.8 | 70.7 | 65.5 | 59.0 | 35.7 | 24.5 | ||
72.0 | 72.1 | 81.1 | 75.9 | 47.1 | 48.3 | 60.3 | 35.5 | 61.2 | 58.4 | 36.2 | 36.0 | ||
Llama 33B | temp with | 80.5 | 75.0 | 95.2 | 93.3 | 65.2 | 66.7 | 75.2 | 73.5 | 79.0 | 77.1 | 80.0 | 70.7 |
78.7 | 71.5 | 95.2 | 92.8 | 68.1 | 67.7 | 75.1 | 73.8 | 77.4 | 75.4 | 73.3 | 62.3 | ||
69.2 | 69.5 | 93.9 | 92.9 | 62.1 | 66.2 | 71.3 | 70.1 | 72.8 | 70.0 | 67.4 | 63.5 |
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||
with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | ||
OpenLlama 3B | temp w/o | 41.5 | 54.6 | 14.3 | 73.2 | 36.5 | 42.0 | 29.4 | 21.7 | 24.7 | 45.7 | 0.7 | 3.5 |
42.2 | 52.2 | 18.5 | 79.9 | 39.7 | 42.2 | 22.6 | 22.5 | 49.8 | 57.1 | 3.1 | 6.5 | ||
36.3 | 35.5 | 83.0 | 83.4 | 43.2 | 41.9 | 16.3 | 18.7 | 54.4 | 56.8 | 1.2 | 4.2 | ||
Llama 7B | temp w/o | 50.4 | 56.6 | 68.2 | 56.1 | 55.3 | 48.5 | 0.2 | 1.3 | 40.7 | 42.5 | 28.5 | 18.8 |
46.1 | 50.6 | 61.9 | 55.1 | 43.5 | 44.7 | 0.0 | 0.2 | 43.7 | 49.9 | 27.1 | 26.5 | ||
21.4 | 12.7 | 86.2 | 66.5 | 54.4 | 55.6 | 0.0 | 0.0 | 56.0 | 55.6 | 39.4 | 35.1 | ||
Llama 13B | temp w/o | 66.9 | 77.0 | 65.6 | 87.9 | 51.8 | 53.1 | 0.1 | 0.1 | 57.5 | 53.7 | 16.7 | 21.9 |
72.9 | 76.6 | 83.0 | 89.5 | 45.5 | 48.4 | 0.0 | 0.0 | 53.6 | 52.8 | 16.0 | 20.1 | ||
79.2 | 77.7 | 77.5 | 47.5 | 56.8 | 43.2 | 0.0 | 0.0 | 54.8 | 53.7 | 4.3 | 2.4 | ||
Llama 33B | temp w/o | 77.3 | 78.2 | 17.3 | 88.9 | 65.4 | 69.3 | 31.0 | 41.7 | 71.8 | 65.8 | 23.8 | 23.0 |
72.9 | 72.4 | 29.2 | 87.4 | 65.6 | 70.9 | 14.9 | 37.9 | 70.6 | 67.8 | 19.9 | 21.1 | ||
69.5 | 74.3 | 92.6 | 92.8 | 70.0 | 70.8 | 42.0 | 20.3 | 67.0 | 61.3 | 23.1 | 18.5 |
Appendix S Significance test for the representation-level ablation
@@
In this section, we report the p-value of all the pair-wise comparisons in the representation-level ablation experiments in Table 2 and Table 3. Results are shown in Table 31. Most of the ablation results show significant difference among different ablation scenarios.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||
P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | ||
OpenLlama 3B | temp <-> cont | 0.0000581 | T | 0.0000000 | T | 0.1952165 | F | 0.0000000 | T | 0.0042427 | T | 0.0027293 | T |
temp <-> stop | 0.0001605 | T | 0.0571278 | F | 0.0242797 | T | 0.0000663 | T | 0.0319815 | T | 0.0985942 | F | |
cont <-> stop | 0.0023957 | T | 0.0000001 | T | 0.1940792 | F | 0.0000000 | T | 0.0000073 | T | 0.0000698 | T | |
temp_cont <-> cont_stop | 0.0000065 | T | 0.0385760 | T | 0.3206221 | F | 0.0000000 | T | 0.1237570 | F | 0.0544049 | F | |
temp_stop <-> cont_stop | 0.0001166 | T | 0.4514005 | F | 0.2549225 | F | 0.0000001 | T | 0.0545474 | F | 0.0534521 | F | |
temp_cont <-> temp_stop | 0.0000507 | T | 0.0096752 | T | 0.0005775 | T | 0.0000000 | T | 0.1208140 | F | 0.3193696 | F | |
Llama 7B | temp <-> cont | 0.0000000 | T | 0.0000001 | T | 0.0000020 | T | 0.0000001 | T | 0.0000004 | T | 0.0000001 | T |
temp <-> stop | 0.0000083 | T | 0.0101283 | T | 0.0002883 | T | 0.1438193 | F | 0.0000000 | T | 0.0000031 | T | |
cont <-> stop | 0.0000060 | T | 0.0000001 | T | 0.0019529 | T | 0.0000001 | T | 0.3392237 | F | 0.0016487 | T | |
temp_cont <-> cont_stop | 0.0000115 | T | 0.0030175 | T | 0.0005950 | T | 0.0000000 | T | 0.0000000 | T | 0.0001649 | T | |
temp_stop <-> cont_stop | 0.0002004 | T | 0.0227328 | T | 0.0015468 | T | 0.0000000 | T | 0.0000001 | T | 0.0000094 | T | |
temp_cont <-> temp_stop | 0.0086396 | T | 0.0003632 | T | 0.0089932 | T | 0.0000007 | T | 0.1637608 | F | 0.1553081 | F | |
Llama 13B | temp <-> cont | 0.0000000 | T | 0.0000000 | T | 0.0000082 | T | 0.0000001 | T | 0.0006445 | T | 0.1060226 | F |
temp <-> stop | 0.0003841 | T | 0.0000012 | T | 0.0034370 | T | 0.0002018 | T | 0.0000000 | T | 0.0010178 | T | |
cont <-> stop | 0.0000000 | T | 0.0000202 | T | 0.0002820 | T | 0.0000000 | T | 0.0098209 | T | 0.0022848 | T | |
temp_cont <-> cont_stop | 0.0010838 | T | 0.0000730 | T | 0.0004557 | T | 0.0000048 | T | 0.0000001 | T | 0.0002364 | T | |
temp_stop <-> cont_stop | 0.0007763 | T | 0.0000310 | T | 0.0016544 | T | 0.0000000 | T | 0.0000000 | T | 0.0000888 | T | |
temp_cont <-> temp_stop | 0.4411518 | F | 0.1158895 | F | 0.3323328 | F | 0.0000000 | T | 0.3148253 | F | 0.0144961 | T | |
Llama 33B | temp <-> cont | 0.0000000 | T | 0.0000003 | T | 0.1534319 | F | 0.0000000 | T | 0.0000000 | T | 0.0002244 | T |
temp <-> stop | 0.0007359 | T | 0.0048547 | T | 0.1797405 | F | 0.0000023 | T | 0.0000002 | T | 0.0008789 | T | |
cont <-> stop | 0.0000000 | T | 0.0000003 | T | 0.0204911 | T | 0.0000000 | T | 0.0002626 | T | 0.4319440 | F | |
temp_cont <-> cont_stop | 0.0001365 | T | 0.0788756 | F | 0.0032131 | T | 0.0000000 | T | 0.0000098 | T | 0.0003242 | T | |
temp_stop <-> cont_stop | 0.0006045 | T | 0.0609501 | F | 0.0165374 | T | 0.0000011 | T | 0.0000009 | T | 0.0003821 | T | |
temp_cont <-> temp_stop | 0.0012936 | T | 0.3583931 | F | 0.1415489 | F | 0.0001034 | T | 0.0009055 | T | 0.3979685 | F |
Appendix T Template used for the random string experiments
@@
Appendix U Supplemental Experiments for the Repetition Characteristic
@@
dog: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
@@
bunny: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe … |
@@
dog: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
@@
dog: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe… |
@@
article: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
@@
sentence: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe… |
@@
article: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
@@
article: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe… |
@@
Models | Setting | AGNews | DBPedia | TREC | Avg. |
OpenLlama 3B | Template 1 | 33.56 | 9.72 | 5.84 | 16.37 |
Template 2 | 55.56 | 61.44 | 22.36 | 46.45 | |
Template 3 | 48.24 | 50.16 | 24.84 | 41.08 | |
Template 4 | 69.64 | 63.20 | 20.96 | 51.27 | |
Llama 7B | Template 1 | 6.00 | 0.24 | 12.92 | 6.39 |
Template 2 | 26.52 | 51.20 | 25.32 | 34.35 | |
Template 3 | 19.80 | 62.28 | 1.88 | 27.99 | |
Template 4 | 36.44 | 64.68 | 18.68 | 39.93 | |
Llama 13B | Template 1 | 7.40 | 0.00 | 5.68 | 4.36 |
Template 2 | 46.68 | 71.40 | 35.80 | 51.29 | |
Template 3 | 15.60 | 76.36 | 4.96 | 32.31 | |
Template 4 | 49.08 | 75.56 | 23.80 | 49.48 | |
Llama 2 7B | Template 1 | 52.84 | 12.08 | 35.60 | 33.51 |
Template 2 | 75.60 | 82.80 | 56.04 | 71.48 | |
Template 3 | 32.96 | 76.56 | 7.04 | 38.85 | |
Template 4 | 70.16 | 80.96 | 58.24 | 69.79 | |
Llama 2 13B | Template 1 | 8.28 | 0.04 | 2.68 | 3.67 |
Template 2 | 19.80 | 44.52 | 13.92 | 26.08 | |
Template 3 | 5.84 | 59.88 | 1.72 | 22.48 | |
Template 4 | 28.60 | 65.04 | 12.64 | 35.43 | |
Mistral 7B | Template 1 | 27.68 | 0.20 | 17.96 | 15.28 |
Template 2 | 67.48 | 67.60 | 31.20 | 55.43 | |
Template 3 | 2.64 | 47.68 | 4.04 | 18.12 | |
@@
Appendix V Supplemental Experiments for the Structural Cue Characteristic
In this section, we describe a set of supplemental experiments, which support the characteristic of structural cues from the perspective of representation-level ablation. An intuitive method to verify the effect of the structural cues would be using the same random strings to replace and , making it harder for a model to parse the structure of the text. However, this would bring the factor of repetition into the process, potentially confounding the results. Hence, we instead design a one-shot Randomfixed experiment. The one-shot Randomfixed setting allows us to control both the characteristics of lexical meaning and repetition since the templates are made up of random strings and there is only one training demonstration. With these two characteristics controlled, we use the masking ablation method from Section 5.2.1 to confirm to what extent these random string tokens can function effectively as delimiters between inputs and outputs in ICL prompts. Specifically, we include results from the Zero-shot temp and Zero-shot “:” scenarios, as well as the standard results of one-shot Randomfixed, for a more comprehensive analysis, shown in Table 35. Examples of all the different model variants are shown in Appendix W.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | One-shot Randomfixed | 47.5 | 51.8 | 32.6 | 19.4 | 51.8 | 42.4 | 40.9 |
Zero-shottemp | 39.5 | 49.8 | 27.7 | 13.3 | 49.8 | 44.9 | 37.5 | |
Zero-shot“:” | 31.5 | 35.9 | 23.8 | 8.0 | 35.9 | 33.8 | 28.2 | |
Llama 7B | One-shot Randomfixed | 3.9 | 16.9 | 3.5 | 9.6 | 16.9 | 10.4 | 10.2 |
Zero-shottemp | 2.1 | 15.5 | 7.6 | 3.7 | 15.5 | 5.4 | 8.3 | |
Zero-shot“:” | 3.6 | 7.5 | 14.6 | 3.0 | 7.5 | 6.8 | 7.2 | |
Llama 13B | One-shot Randomfixed | 46.1 | 47.5 | 25.0 | 50.8 | 47.5 | 21.4 | 39.7 |
Zero-shottemp | 29.2 | 48.9 | 36.1 | 35.7 | 48.9 | 14.0 | 35.5 | |
Zero-shot“:” | 14.3 | 22.4 | 25.4 | 22.5 | 22.4 | 28.9 | 22.7 | |
Llama 33B | One-shot Randomfixed | 69.7 | 53.0 | 37.8 | 72.8 | 53.0 | 37.6 | 54.0 |
Zero-shottemp | 61.2 | 56.3 | 41.1 | 69.2 | 56.3 | 43.0 | 54.5 | |
@@ Appendix V Supplemental Experiments for the Structural Cue Characteristic
Appendix W Discussion about the Characteristic of Structural Cue

### v4

@@ Appendix I Stopword Tokens
The full results on all the six datasets are shown in Table 32 and Table 33. Most of the results align with our descriptions in Section D, where the performance-critical tokens should be utilized together to provide the best performance and that removing some of them would cause performance degeneration, demonstrated by the performance decrease from Table 32, or instability issues, shown by Table 33.
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||||
| with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | ||||
temp with | 56.5 | 47.4 | 86.7 | 83.7 | 27.1 | 26.5 | 62.2 | 59.8 | 56.4 | 56.0 | 52.3 | 56.1 | ||
| 50.3 | 47.1 | 85.7 | 84.4 | 28.9 | 24.4 | 57.7 | 57.7 | 56.5 | 56.1 | 53.2 | 55.2 | ||||
| 34.6 | 32.7 | 86.9 | 82.3 | 28.2 | 31.2 | 55.5 | 54.1 | 58.3 | 59.2 | 55.4 | 58.3 | ||||
temp with | 70.8 | 57.3 | 90.2 | 87.1 | 58.4 | 46.7 | 66.2 | 63.8 | 66.3 | 59.5 | 73.5 | 69.6 | ||
| 62.7 | 55.1 | 91.6 | 87.1 | 52.8 | 43.3 | 61.6 | 61.8 | 58.9 | 56.5 | 59.2 | 55.7 | ||||
| 50.8 | 48.6 | 84.9 | 82.8 | 46.0 | 50.2 | 57.9 | 55.2 | 56.7 | 56.7 | 66.2 | 64.5 | ||||
temp with | 80.0 | 76.2 | 92.3 | 89.1 | 58.6 | 54.0 | 76.9 | 71.4 | 68.5 | 59.8 | 47.7 | 35.0 | ||
| 79.9 | 76.3 | 91.5 | 88.9 | 55.1 | 47.8 | 75.8 | 70.7 | 65.5 | 59.0 | 35.7 | 24.5 | ||||
| 72.0 | 72.1 | 81.1 | 75.9 | 47.1 | 48.3 | 60.3 | 35.5 | 61.2 | 58.4 | 36.2 | 36.0 | ||||
temp with | 80.5 | 75.0 | 95.2 | 93.3 | 65.2 | 66.7 | 75.2 | 73.5 | 79.0 | 77.1 | 80.0 | 70.7 | ||
| 78.7 | 71.5 | 95.2 | 92.8 | 68.1 | 67.7 | 75.1 | 73.8 | 77.4 | 75.4 | 73.3 | 62.3 | ||||
| 69.2 | 69.5 | 93.9 | 92.9 | 62.1 | 66.2 | 71.3 | 70.1 | 72.8 | 70.0 | 67.4 | 63.5 | ||||
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||||
| with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | ||||
temp w/o | 41.5 | 54.6 | 14.3 | 73.2 | 36.5 | 42.0 | 29.4 | 21.7 | 24.7 | 45.7 | 0.7 | 3.5 | ||
| 42.2 | 52.2 | 18.5 | 79.9 | 39.7 | 42.2 | 22.6 | 22.5 | 49.8 | 57.1 | 3.1 | 6.5 | ||||
| 36.3 | 35.5 | 83.0 | 83.4 | 43.2 | 41.9 | 16.3 | 18.7 | 54.4 | 56.8 | 1.2 | 4.2 | ||||
temp w/o | 50.4 | 56.6 | 68.2 | 56.1 | 55.3 | 48.5 | 0.2 | 1.3 | 40.7 | 42.5 | 28.5 | 18.8 | ||
| 46.1 | 50.6 | 61.9 | 55.1 | 43.5 | 44.7 | 0.0 | 0.2 | 43.7 | 49.9 | 27.1 | 26.5 | ||||
| 21.4 | 12.7 | 86.2 | 66.5 | 54.4 | 55.6 | 0.0 | 0.0 | 56.0 | 55.6 | 39.4 | 35.1 | ||||
temp w/o | 66.9 | 77.0 | 65.6 | 87.9 | 51.8 | 53.1 | 0.1 | 0.1 | 57.5 | 53.7 | 16.7 | 21.9 | ||
| 72.9 | 76.6 | 83.0 | 89.5 | 45.5 | 48.4 | 0.0 | 0.0 | 53.6 | 52.8 | 16.0 | 20.1 | ||||
| 79.2 | 77.7 | 77.5 | 47.5 | 56.8 | 43.2 | 0.0 | 0.0 | 54.8 | 53.7 | 4.3 | 2.4 | ||||
temp w/o | 77.3 | 78.2 | 17.3 | 88.9 | 65.4 | 69.3 | 31.0 | 41.7 | 71.8 | 65.8 | 23.8 | 23.0 | ||
| 72.9 | 72.4 | 29.2 | 87.4 | 65.6 | 70.9 | 14.9 | 37.9 | 70.6 | 67.8 | 19.9 | 21.1 | ||||
| 69.5 | 74.3 | 92.6 | 92.8 | 70.0 | 70.8 | 42.0 | 20.3 | 67.0 | 61.3 | 23.1 | 18.5 | ||||
Appendix T Significance test for the representation-level ablation
@@
In this section, we report the p-value of all the pair-wise comparisons in the representation-level ablation experiments in Table 3. Results are shown in Table 34. Most of the ablation results show significant difference among different ablation scenarios.
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||||
| P-value | p ¡ 0.05 | P-value | p ¡ 0.05 | P-value | p ¡ 0.05 | P-value | p ¡ 0.05 | P-value | p ¡ 0.05 | P-value | p ¡ 0.05 | ||||
temp (-) cont | 0.0000581 | T | 0.0000000 | T | 0.1952165 | F | 0.0000000 | T | 0.0042427 | T | 0.0027293 | T | ||
| temp (-) stop | 0.0001605 | T | 0.0571278 | F | 0.0242797 | T | 0.0000663 | T | 0.0319815 | T | 0.0985942 | F | |||
| cont (-) stop | 0.0023957 | T | 0.0000001 | T | 0.1940792 | F | 0.0000000 | T | 0.0000073 | T | 0.0000698 | T | |||
| temp_cont (-) cont_stop | 0.0000065 | T | 0.0385760 | T | 0.3206221 | F | 0.0000000 | T | 0.1237570 | F | 0.0544049 | F | |||
| temp_stop (-) cont_stop | 0.0001166 | T | 0.4514005 | F | 0.2549225 | F | 0.0000001 | T | 0.0545474 | F | 0.0534521 | F | |||
| temp_cont (-) temp_stop | 0.0000507 | T | 0.0096752 | T | 0.0005775 | T | 0.0000000 | T | 0.1208140 | F | 0.3193696 | F | |||
temp (-) cont | 0.0000000 | T | 0.0000001 | T | 0.0000020 | T | 0.0000001 | T | 0.0000004 | T | 0.0000001 | T | ||
| temp (-) stop | 0.0000083 | T | 0.0101283 | T | 0.0002883 | T | 0.1438193 | F | 0.0000000 | T | 0.0000031 | T | |||
| cont (-) stop | 0.0000060 | T | 0.0000001 | T | 0.0019529 | T | 0.0000001 | T | 0.3392237 | F | 0.0016487 | T | |||
| temp_cont (-) cont_stop | 0.0000115 | T | 0.0030175 | T | 0.0005950 | T | 0.0000000 | T | 0.0000000 | T | 0.0001649 | T | |||
| temp_stop (-) cont_stop | 0.0002004 | T | 0.0227328 | T | 0.0015468 | T | 0.0000000 | T | 0.0000001 | T | 0.0000094 | T | |||
| temp_cont (-) temp_stop | 0.0086396 | T | 0.0003632 | T | 0.0089932 | T | 0.0000007 | T | 0.1637608 | F | 0.1553081 | F | |||
temp (-) cont | 0.0000000 | T | 0.0000000 | T | 0.0000082 | T | 0.0000001 | T | 0.0006445 | T | 0.1060226 | F | ||
| temp (-) stop | 0.0003841 | T | 0.0000012 | T | 0.0034370 | T | 0.0002018 | T | 0.0000000 | T | 0.0010178 | T | |||
| cont (-) stop | 0.0000000 | T | 0.0000202 | T | 0.0002820 | T | 0.0000000 | T | 0.0098209 | T | 0.0022848 | T | |||
| temp_cont (-) cont_stop | 0.0010838 | T | 0.0000730 | T | 0.0004557 | T | 0.0000048 | T | 0.0000001 | T | 0.0002364 | T | |||
| temp_stop (-) cont_stop | 0.0007763 | T | 0.0000310 | T | 0.0016544 | T | 0.0000000 | T | 0.0000000 | T | 0.0000888 | T | |||
| temp_cont (-) temp_stop | 0.4411518 | F | 0.1158895 | F | 0.3323328 | F | 0.0000000 | T | 0.3148253 | F | 0.0144961 | T | |||
temp (-) cont | 0.0000000 | T | 0.0000003 | T | 0.1534319 | F | 0.0000000 | T | 0.0000000 | T | 0.0002244 | T | ||
| temp (-) stop | 0.0007359 | T | 0.0048547 | T | 0.1797405 | F | 0.0000023 | T | 0.0000002 | T | 0.0008789 | T | |||
| cont (-) stop | 0.0000000 | T | 0.0000003 | T | 0.0204911 | T | 0.0000000 | T | 0.0002626 | T | 0.4319440 | F | |||
| temp_cont (-) cont_stop | 0.0001365 | T | 0.0788756 | F | 0.0032131 | T | 0.0000000 | T | 0.0000098 | T | 0.0003242 | T | |||
| temp_stop (-) cont_stop | 0.0006045 | T | 0.0609501 | F | 0.0165374 | T | 0.0000011 | T | 0.0000009 | T | 0.0003821 | T | |||
| temp_cont (-) temp_stop | 0.0012936 | T | 0.3583931 | F | 0.1415489 | F | 0.0001034 | T | 0.0009055 | T | 0.3979685 | F | |||
Appendix U Template used for the random string experiments
@@
Appendix V Supplemental Experiments for the Repetition Characteristic
@@
| dog: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying ”thousands” of fee-paying astronauts could be sent into orbit in the near future. |
@@
| bunny: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: ”Europe … |
@@
| dog: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying ”thousands” of fee-paying astronauts could be sent into orbit in the near future. |
@@
| dog: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: ”Europe… |
@@
| article: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying ”thousands” of fee-paying astronauts could be sent into orbit in the near future. |
@@
| sentence: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: ”Europe… |
@@
| article: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying ”thousands” of fee-paying astronauts could be sent into orbit in the near future. |
@@
| article: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: ”Europe… |
@@
| Models | Setting | AGNews | DBPedia | TREC | Avg. | ||
Template 1 | 33.56 | 9.72 | 5.84 | 16.37 | ||
| Template 2 | 55.56 | 61.44 | 22.36 | 46.45 | |||
| Template 3 | 48.24 | 50.16 | 24.84 | 41.08 | |||
| Template 4 | 69.64 | 63.20 | 20.96 | 51.27 | |||
Template 1 | 6.00 | 0.24 | 12.92 | 6.39 | ||
| Template 2 | 26.52 | 51.20 | 25.32 | 34.35 | |||
| Template 3 | 19.80 | 62.28 | 1.88 | 27.99 | |||
| Template 4 | 36.44 | 64.68 | 18.68 | 39.93 | |||
Template 1 | 7.40 | 0.00 | 5.68 | 4.36 | ||
| Template 2 | 46.68 | 71.40 | 35.80 | 51.29 | |||
| Template 3 | 15.60 | 76.36 | 4.96 | 32.31 | |||
| Template 4 | 49.08 | 75.56 | 23.80 | 49.48 | |||
Template 1 | 52.84 | 12.08 | 35.60 | 33.51 | ||
| Template 2 | 75.60 | 82.80 | 56.04 | 71.48 | |||
| Template 3 | 32.96 | 76.56 | 7.04 | 38.85 | |||
| Template 4 | 70.16 | 80.96 | 58.24 | 69.79 | |||
Template 1 | 8.28 | 0.04 | 2.68 | 3.67 | ||
| Template 2 | 19.80 | 44.52 | 13.92 | 26.08 | |||
| Template 3 | 5.84 | 59.88 | 1.72 | 22.48 | |||
| Template 4 | 28.60 | 65.04 | 12.64 | 35.43 | |||
Template 1 | 27.68 | 0.20 | 17.96 | 15.28 | ||
| Template 2 | 67.48 | 67.60 | 31.20 | 55.43 | |||
| Template 3 | 2.64 | 47.68 | 4.04 | 18.12 | |||
@@
Appendix W Supplemental Experiments for the Structural Cue Characteristic
In this section, we describe a set of supplemental experiments, which support the characteristic of structural cues from the perspective of representation-level ablation. An intuitive method to verify the effect of the structural cues would be using the same random strings to replace and , making it harder for a model to parse the structure of the text. However, this would bring the factor of repetition into the process, potentially confounding the results. Hence, we instead design a one-shot Randomfixed experiment. The one-shot Randomfixed setting allows us to control both the characteristics of lexical meaning and repetition since the templates are made up of random strings and there is only one training demonstration. With these two characteristics controlled, we use the masking ablation method from Section 5.2 to confirm to what extent these random string tokens can function effectively as delimiters between inputs and outputs in ICL prompts. Specifically, we include results from the Zero-shot temp and Zero-shot “:” scenarios, as well as the standard results of one-shot Randomfixed, for a more comprehensive analysis, shown in Table 38. Examples of all the different model variants are shown in Appendix X.
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
One-shot Randomfixed | 47.5 | 51.8 | 32.6 | 19.4 | 51.8 | 42.4 | 40.9 | ||
| Zero-shottemp | 39.5 | 49.8 | 27.7 | 13.3 | 49.8 | 44.9 | 37.5 | |||
| Zero-shot“:” | 31.5 | 35.9 | 23.8 | 8.0 | 35.9 | 33.8 | 28.2 | |||
One-shot Randomfixed | 3.9 | 16.9 | 3.5 | 9.6 | 16.9 | 10.4 | 10.2 | ||
| Zero-shottemp | 2.1 | 15.5 | 7.6 | 3.7 | 15.5 | 5.4 | 8.3 | |||
| Zero-shot“:” | 3.6 | 7.5 | 14.6 | 3.0 | 7.5 | 6.8 | 7.2 | |||
One-shot Randomfixed | 46.1 | 47.5 | 25.0 | 50.8 | 47.5 | 21.4 | 39.7 | ||
| Zero-shottemp | 29.2 | 48.9 | 36.1 | 35.7 | 48.9 | 14.0 | 35.5 | |||
| Zero-shot“:” | 14.3 | 22.4 | 25.4 | 22.5 | 22.4 | 28.9 | 22.7 | |||
One-shot Randomfixed | 69.7 | 53.0 | 37.8 | 72.8 | 53.0 | 37.6 | 54.0 | ||
| Zero-shottemp | 61.2 | 56.3 | 41.1 | 69.2 | 56.3 | 43.0 | 54.5 | |||
@@
Appendix X Discussion about the Characteristic of Structural Cue